# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented Large Language Models

### v4

@@ 5. Evaluation
In this section, we conduct an extensive evaluation of TPIA, demonstrating its threat to Code LLMs from the following perspectives:
@@
    TPIA can successfully achieve representative malicious objectives on different open-source Code LLMs with high effectiveness and stealthiness. (Section 5.2)
@@
    TPIA has a sufficient transferability to confirm its feasibility in real-world commercial applications under gray-box and black-box settings. (Section 5.3)
@@
    TPIA’s modules are all effective in improving attack performance. (Section 5.4)
5.1. Experiment Settings
Table 1. The summary of objectives and datasets. We list the programming languages that each objective/dataset is applicable: C/C++ (C), Python (P), Java (J). The last row of the table shows the total number of test data we design. Some objectives apply only to C/C++, resulting in a total of 31 applicable cases.
@@ 5.1. Experiment Settings
Table 2. The ASR across different Code LLMs. The experiment targets four code completion datasets and tests the success rates of various attack methods using three Code LLMs. As shown in Table 1, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets. Best attack results are marked in bold.
@@
Malicious Objectives and Code Datasets. We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) (Top, 2024) across 3 programming languages. The straightforward objectives include using MODE_ECB for encryption (dubbed ‘ST0’), using ssl3 protocol (dubbed ‘ST1’), and executing “rm -rf /” (dubbed ‘ST2’). ST0 and ST1 are studied in classical code backdoor attacks (Schuster et al., 2021), while ST2 is a famous destructive instruction. CWE vulnerabilities are generally selected for experiments in code model research (He and Vechev, 2023; Yan et al., 2024), and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA). We consider three mainstream programming languages: Python, Java, and C/C++ to design the threat case for each objective. All the cases are constructed by the method described in Section 4.2 based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2). We will discuss the cases in detail in Appendix A.2. As shown in Table 1, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases. In addition, we collect corresponding code completion datasets, including Humaneval (Chen et al., 2021), Humaneval-x (Zheng et al., 2023), CodeXGLUE (Lu et al., 2021), MBPP (Austin et al., 2021), and Eval-Plus (Liu et al., 2024). The items in these datasets contain prompts and corresponding outputs across different code completion tasks, which can simulate unpredictable users’ instructions and context in real-world scenarios. Currently, most code completion datasets only cover Python, as shown in Table 1. As a result, we have 7 sub-datasets (i.e., the subsection of a dataset containing a particular programming language) containing more than 14,000 data items totally.
@@
Model Selection. In our experiment, we select three models from two representative Code LLM series, including CodeGemma (Team et al., 2024) and Code-GeeX (Zheng et al., 2023), for evaluation. CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama (Roziere et al., 2023) on code completion tasks. We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales. CodeGeeX series have corresponding applications (cod, 2024) with more than 535k downloads (on VScode (VSC, 2024) Apps shop). We select CodeGeeX2-6b, a version with stable performance. We exclude earlier small-scale models, such as CodeBERT (Feng et al., 2020) and CodeT5 (Wang et al., 2021), because they perform much worse than large-scale models in code completion tasks.
@@
Baseline Selection. We compare our TPIA with GCG-based and manually designed IPI attacks. In particular, the vanilla GCG (Zou et al., 2023) is a representative PI attack and does not meet the requirements of our attack scenarios. Therefore, we adapt GCG into an IPI attack that is applicable to Code LLMs, dubbed ‘C-GCG’ (code-oriented GCG). The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work (Greshake et al., 2023). IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments (as shown in Figure 11). Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have some inherent insecurity factors. In particular, we exclude backdoor attacks and adversarial attacks as our baselines because it is difficult to fairly compare these methods with TPIA. Specifically, backdoor attacks require control over the model training process, which is an overly strong assumption for our threat model. Adversarial attacks, on the other hand, cannot achieve target-specific attacks, leading to unreliable threat levels. We provide more explanations in the experiments in Appendix A.3.
@@
Evaluation Metrics. In this paper, we use attack success rate (ASR) to evaluate the effectiveness, not bad cases rate (NBR) and selective threat factor (STF) to evaluate the flexibility, and number of injected tokens and chars (NIT and NIC) to evaluate the stealthiness. For ASR, we consider an attack successful only when the generated code has equivalent functionality to the target code and appears in the correct position. In other words, the generated code must not only have malicious functionality but also adhere to the code compilation syntax rules. NBR and STF are new metrics introduced for our target-specific attack scenarios. NBR represents the rate of feasible cases. The more cases remaining after excluding the unusable ones with low ASR (below a threshold $r_{1}$), the greater variety of malicious objectives the attack can achieve; STF represents the rate of successful cases and their threat level. The more cases with sufficiently high ASRs (above a threshold $r_{2}$) and larger overall ASR value, the greater the upper limit of the attack’s threat. Specifically, NBR and STF can be formulated as follows:
(7) NBR $=1-\frac{\text{\# of cases where ASR}<r_{1}}{\text{Total \# of cases}},$
@@
where $r_{1}$ and $r_{2}$ are set as 20% and 80% in our evaluation. More details about the selection of metrics and values are in Appendix A.2.
Hyperparameters and Computational Facilities. We conduct experiments using the following parameter settings to evaluate our approach. During gradient calculation, we utilize five noisy code samples, and the random seed for initializing perturbations is set to 100. For forward reasoning enhancement in Section 4.4, the number of tokens reinforced is set to $h=2$, as defined in Equation 6. For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate ($r$ in Equation 6) is 0.4, the number of keywords is 2, and the perturbation length is 10. Tokenization is performed using Gemma’s tokenizer.
We run Algorithm 1 on an NVIDIA RTX A6000 Graphics Card with the above parameter settings, employing a float16 precision format without quantization. In general, the perturbation generation process requires approximately 0.5-1 hour for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b. These computational times ensure the generation of high-quality perturbations tailored to the respective model sizes.
5.2. Evaluation on Open-source Code LLMs
@@
Effectiveness. We hereby evaluate whether TPIA can successfully induce the Code LLMs to generate attacker-specified targets. We control the number of injected tokens for three attack methods in this experiment. For TPIA, we use its default settings: perturbation length of 10 and 2 keywords, while C-GCG uses a prefix length of 12. Since IPI-M needs sufficient tokens to be effective, we do not set a fixed length for this attack but limit the number of used tokens under 30 (details of IPI-M are described in Appendix A.2). Table 2 provides the average ASR for four methods. The effectiveness of TPIA significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
Comparing the three Code LLMs, the CodeGemma-7b model has the highest security in normal conditions (8.4% for No Attack) and is the least likely to generate vulnerable code. However, it is the most susceptible model to TPIA, which may be attributed to its strong contextual learning ability. On the other hand, the worst-performing CodeGemma-2b is more easily affected by simpler and more direct baseline methods such as C-GCG and IPI-M; Comparing the three programming languages, Python is the most susceptible language to TPIA due to its concise syntax and the smallest number of tokens consumed; Besides, TPIA is relatively less effective on CodeXGLUE, which mostly due to its larger data length and more complex data content. In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M. Therefore, we believe that an increase in the contextual complexity can reduce the ASR of TPIA. In other words, when the contextual code becomes complex, the strength of TPIA (e.g., the length of perturbations) needs to be increased accordingly to ensure the attack’s effectiveness. According to our subsequent experiments in Section 5.3 and Section 5.4, increasing the length of TPIA or Top-$k$ value can achieve better attack performance.
An interesting phenomenon is that IPI-M can sometimes enhance the security of the model’s output (producing a lower ASR than ‘No Attack’). We speculate that this is mostly because, in some cases, the generated perturbation contains information on the potential positions where vulnerabilities might occur but does not clearly convey the content of the malicious behavior to Code LLMs. As such, Code LLMs actually strengthen their reasoning about the locations of potential weaknesses and produce more secure outputs. It suggests that (parts of) our TPIA (e.g., those used in IPI-M) could potentially be used for enhancing the security of Code LLMs, by automatically optimizing the unfunctional parts of reference code to improve the security of the generated code. We will further explore it in our future work.
Table 3. The attack evaluation results of NBR and STF against open-source Code LLMs.
@@
As shown in Table 3, TPIA has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible. Additionally, TPIA scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat. STF and NBR are not always positively correlated. For example, for TPIA, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score. This indicates that there are more significant and effective attack examples in C/C++. While attacks on Java may be more comprehensive, they are harder to achieve stable and significant results. Besides, whether it is NBR or STF, TPIA significantly outperforms the baselines.
Stealthiness. We hereby evaluate the trade-off between effectiveness and stealthiness. As we discussed in Section 3.3, our TPIA can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals. Arguably, human inspection is currently the only feasible approach to detect TPIA attacks. Therefore, we measure the stealthiness of TPIA by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
As shown in Figure 6, ASR and STF show a clear increase as NIT increases. When NIT is above 12, ASR and STF completely overlap, indicating that ASR exceeds 80% across all cases. Before reaching 60% ASR, TPIA attacks require very low NIT/NIC values. If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy. When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation). This demonstrates that enhancing TPIA’s performance does not require significantly increasing the length of the perturbation as with existing methods, adding just a few tokens can be highly effective.
@@
Figure 6. The study of stealthiness. We plot the above two graphs using NIT and ASR as the x-axis respectively. We sampled 11 different NITs, namely NIT 0 (representing No Attack) and NIT 2 to 20 (with a sampling interval of 2).
5.3. Evaluation on Commercial Code LLMs
@@
    Black-Box Setting. A more challenging situation is when the attacker knows nothing about the model used by the target application, or the model itself is closed-source (such as CodeX). This is precisely the case with GitHub Copilot. We simulate this scenario by using perturbations generated by CodeGeeX2-6b to attack GitHub Copilot.
Table 4. The evaluation of transferability. We test TPIA’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
@@
Figure 7. The example of a successful attack in a real-world black-box scenario (Github Copilot) across files. The perturbation is inserted into the comment of a related file and influence the suggestion of LLM in main file.
Settings. We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total). Other settings are the same as those described in Section 5.2.
Results. As shown in Table 4, in the gray-box scenario, TPIA still demonstrates great attack performance. Specifically, TPIA achieves ASRs of 80.4% and 93.3% in Python and C/C++, respectively. Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, TPIA also achieves a satisfying attack performance of 90.4% on Python. The ASRs on C/C++ and Java are 67.7% and 74.4%, respectively. In other words, our TPIA still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent. In addition, TPIA achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility. Overall, TPIA achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to real-world commercial applications.
In Figure 7, we present a representative case of ST2 under the black-box scenario. The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function. The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction.
5.4. Ablation Study

### v5

@@ V Experiments
In this section, we conduct an extensive evaluation of our TPIA, demonstrating its realistic threat to Code LLMs from the following three perspectives:
@@
    TPIA can successfully achieve representative malicious objectives on different open-source Code LLMs with high effectiveness and stealthiness. (Section V-B)
@@
    TPIA has a sufficient transferability to confirm its feasibility in real-world commercial applications under gray-box and black-box settings. (Section V-C)
@@
    TPIA’s modules are all effective in improving attack performance. (Section V-D)
V-A Main Settings
TABLE I: The summary of objectives and datasets. We list the programming languages that each objective/dataset is applicable: C/C++ (C), Python (P), Java (J). The last row of the table shows the total number of test data we design. Some objectives apply only to C/C++, resulting in a total of 31 applicable cases.
@@ V-A Main Settings
TABLE II: The ASR across different Code LLMs. The experiment targets four code completion datasets and tests the success rates of various attack methods using three Code LLMs. As shown in Table I, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets. Best attack results are marked in bold.
@@
Malicious Objectives and Code Datasets. We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages: Python, Java, and C/C++. The straightforward objectives include using MODE_ECB for encryption (dubbed ‘ST0’), using ssl3 protocol (dubbed ‘ST1’), and executing “rm -rf /” (dubbed ‘ST2’). ST0 and ST1 are studied in classical code backdoor attacks [46], while ST2 is a famous destructive instruction. CWE vulnerabilities are generally selected for experiments in code model research [23, 56], and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA). We construct cases using the method described in Section IV-B based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2). As shown in Table I, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases. In addition, we collect corresponding code completion datasets, including Humaneval [14], Humaneval-x [61], CodeXGLUE [38], MBPP [10], and Eval-Plus [35]. The items in these datasets contain prompts and corresponding outputs across different code completion tasks, which can simulate unpredictable users’ instructions and context in real-world scenarios. Some datasets do not cover Java or C/C++, as shown in Table I. As a result, we have 7 sub-datasets (i.e., the subsection of a dataset containing a particular programming language) containing more than 14,000 data items totally.
@@
Model Selection. In our experiment, we select three models from two representative Code LLM series, including CodeGemma [49] and Code-GeeX [61], for evaluation. CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama [44] on code completion tasks. We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales. CodeGeeX series have corresponding applications [2] with more than 535k downloads (on VScode [6] Apps shop). We select CodeGeeX2-6b, a version with stable performance. We exclude earlier small-scale code-oriented models, such as CodeBERT [18] and CodeT5 [51], because large-scale code-oriented models perform much better than them in code completion tasks.
@@
Baseline Selection. We compare our TPIA with GCG-based and manually designed IPI attacks. In particular, the vanilla GCG [63] is a representative PI attack and does not meet the requirements of our attack scenarios. Therefore, we adapt GCG into an IPI attack that is applicable to Code LLMs, dubbed ‘C-GCG’ (code-oriented GCG). The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work [22]. IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments. Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have inherent insecurity factor.
@@
Evaluation Metrics. In this paper, we use ASR to evaluate the effectiveness, not bad cases rate (NBR) and selective threat factor (STF) to evaluate the flexibility, and number of injected tokens and chars (NIT and NIC) to evaluate the stealthiness. For ASR, we consider an attack successful only when the generated code has equivalent functionality to the target code and appears in the correct position. In other words, the generated code must not only have malicious functionality but also adhere to the code compilation syntax rules. NBR and STF are new metrics introduced for our target-specific attack scenarios. NBR represents the rate of feasible cases. The more cases remaining after excluding the unusable ones with low ASR (below a threshold $r_{1}$), the greater variety of malicious objectives the attack can achieve; STF represents the rate of successful cases and their threat level. The more cases with sufficiently high ASRs (above a threshold $r_{2}$) and larger overall ASR value, the greater the upper limit of the attack’s threat. Their formulations are as follows:
NBR $=1-\frac{\text{\# of cases where ASR}<r_{1}}{\text{Total \# of cases}},$ (8)
@@
where $r_{1}$ and $r_{2}$ are set as 20% and 80% in our evaluation.
Hyperparameters and Computational Facilities. We conduct experiments using the following parameter settings to evaluate our approach. During gradient calculation, we utilize five noisy code samples, and the random seed for initializing perturbations is set to 100. For forward reasoning enhancement in Section IV-D, the number of tokens reinforced is set to $h=2$, as defined in Equation 6. For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate ($r$ in Equation 6) is 0.4, the number of keywords is 2, and the perturbation length is 10. Tokenization is performed using Gemma’s tokenizer.
We run Algorithm 1 on an NVIDIA RTX A6000 Graphics Card with the above parameter settings, employing a float16 precision format without quantization. In general, the perturbation generation process requires approximately 0.5-1 hour for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b. These computational times ensure the generation of high-quality perturbations tailored to the respective model sizes across all the experiment threat cases.
V-B Evaluation on Open-source Code LLMs
@@
Effectiveness. We hereby evaluate whether TPIA can successfully induce the Code LLMs to generate attacker-specified targets. We control the number of injected tokens for three attack methods in this experiment. For TPIA, we use its default settings: perturbation length of 10 and 2 keywords, while C-GCG uses a prefix length of 12. Since IPI-M requires a sufficient number of tokens to be effective, we just limit its perturbation to fewer than 30 tokens. Table II provides the average ASR for four methods. The effectiveness of TPIA significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
Comparing the three Code LLMs, the CodeGemma-7b model has the highest security in normal conditions (8.4% for No Attack) and is the least likely to generate vulnerable code. However, it is the most susceptible model to TPIA, which may be attributed to its strong contextual learning ability. Comparing the three programming languages, Python is the most susceptible language to TPIA due to its concise syntax and the smallest number of tokens consumed; Besides, TPIA is relatively less effective on CodeXGLUE, which mostly due to its larger data length and more complex data content. In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M. Therefore, we believe that an increase in the contextual complexity can reduce the ASR of TPIA. In other words, when the contextual code becomes complex, the strength of TPIA (e.g., the length of perturbations) needs to be increased accordingly to ensure the attack’s effectiveness. According to our subsequent experiments in Section V-C and Section V-D, increasing the length of TPIA or the Top-$k$ value can effectively enhance the attack performance of TPIA.
TABLE III: The attack evaluation results of NBR and STF against open-source Code LLMs.
@@
As shown in Table III, TPIA has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible. Additionally, TPIA scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat. STF and NBR are not always positively correlated. For example, for TPIA, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score. This indicates that there are more significant and effective attack examples in C/C++. While attacks on Java may be more comprehensive, they are harder to achieve stable and significant results. Besides, whether it is NBR or STF, TPIA significantly outperforms the baselines.
Stealthiness. We hereby evaluate the trade-off between effectiveness and stealthiness. As we discussed in Section III-C, our TPIA can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals. Arguably, human inspection is currently the only feasible approach to detect TPIA attacks. Therefore, we measure the stealthiness of TPIA by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
As shown in Figure 6, ASR and STF show a clear increase as NIT increases. Before reaching 60% ASR, TPIA attacks require very low NIT/NIC values. If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy. When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation). This demonstrates that enhancing TPIA’s performance does not require significantly increasing the length of the perturbation as with existing methods [22], adding just a few perturbation tokens can be highly effective.
@@
Figure 6: The study of stealthiness. We plot the above two graphs using NIT and ASR as the x-axis respectively. We sampled 11 different NITs, NIT 0 (representing No Attack) and NIT 2 to 20 (with a sampling interval of 2).
V-C Evaluation on Commercial Code LLMs
@@
    Black-Box Setting. A more challenging situation is when the attacker knows nothing about the model used by the target application, or the model itself is closed-source (such as CodeX). We simulate this scenario by attacking GitHub Copilot (supported by CodeX). To align with gray-box setting, we adopt CodeGeeX2-6b to generate malicious perturbations.
TABLE IV: The evaluation of transferability. We test TPIA’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
@@
Figure 7: The example of a successful attack in a real-world black-box scenario (Github Copilot) across files. The perturbation is inserted into the comment of a function. When the user calls a function in a code context, Copilot will be induced to generate malicious code at the specified position (after TRG).
Settings. We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total). Other settings are the same as those described in Section V-B.
Results. As shown in Table IV, in the gray-box scenario, TPIA still demonstrates great attack performance. Specifically, TPIA achieves ASRs of 80.4% and 93.3% in Python and C/C++, respectively. Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, TPIA also achieves a satisfying attack performance of 90.4% on Python. The ASRs on C/C++ and Java are 67.7% and 74.4%, respectively. In other words, our TPIA still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent. In addition, TPIA achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility. Overall, TPIA achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to the commercial applications in the real-world.
In Figure 7, we present a representative case of ST2 under the black-box scenario. The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function. The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction that can cause dangerous results.
V-D Ablation Study